Given the demand function for computer as p=2400-0.5Q determine the coefficient of point of elasticity of demand when p=1800


1
Expert's answer
2021-11-08T20:07:56-0500

"P=2400 -0.4Q \\

0.4Q = 2400 -P \\

Q = \frac{2400}{0.4} - \frac{P}{0.4} \\

Q = 6000 -2.5P \\

\frac{∂Q}{∂p} = \frac{∂}{∂P}(6000 -2.5P) \\

\frac{∂Q}{∂p} = -2.5"

For P=1800

"Q= 6000 -2.5 \times 1800 = 6000 -4500 = 1500"

The coefficient of point of elasticity of demand can be obtained by using the formula:

"e_p = \frac{∂Q}{∂p} \times \frac{P_0}{Q_0} \\

e_p = -2.5 \times \frac{1800}{1500} \\

e_p= -3"